what is the minimum or maximum value of the quadratic formula
f(x)=2x^2-16x+20

Answer by nerdybill(7384)

f(x)=2x^2-16x+20
.
Since the coefficient associated with the x^2 is POSITIVE, we conclude that it is a parabola that opens upwards. Therefore, the vertex will give the MINIMUM.
.
x-value of vertex:
x = -b/(2a)
x = -(-16)/(2*2)
x = (16)/(4)
x = 4
.
y-value of vertex:
f(4)=2*4^2-16(4)+20
f(4)=2*16-16(4)+20
f(4)=32-64+20
f(4)=-32+20
f(4)=-12 (this is the answer they're looking for) (minimum of quadratic)